slope of a tan line to a curve at P (3,3) is 4. Use the equation of this tanngent to find its value for x=9.
Answer by stanbon(75887)

slope of a tan line to a curve at P (3,3) is 4. Use the equation of this tangent to find its value for x=9.
----
Find the equation of the tangent line:
slope = 4
Form = y = mx+b
3 = 4*3+b
b = -9
Equation of the tangent: y = 4x-9
This is the definite derivative of the function "P"
at x = 3
----
Integrate to get P:
P(x) = 4(x^2/2) - 9x
P(x) = 2x^2 -9x
---------------------
P(9) = 2(9^2)-9*9
P(9) = 81
=================
Cheers,
Stan H.